# A Survey on LLM Mid-training

### v1

@@ 2 Overview
Mid-training occupies a distinct functional niche in the training pipeline, situated between general pre-training and specialized post-training stages (Figure 1). Contemporary LLM development frameworks position pre-training as the bedrock stage, where models acquire broad competencies through next-token prediction on web-scale corpora. Emerging as an evolutionary bridge, mid-training preserves these foundational capabilities with the same next-token prediction objective, while systematically enhancing specialized skills through curated data blends—often incorporating high-quality domain-specific content and instruction data. Unlike pre-training, mid-training demonstrates steeper performance gains with higher data efficiency, frequently employing multiple sub-phases for different capability upgrades. In contrast, post-training adopts specialized objectives for SFT/RL and task-specific alignment datasets.
@@ 3.2 Data Synthesis
Distillation leverages powerful LLMs to directly generate data through carefully engineered prompts or to distill smaller synthesis models. This approach is heavily reliant on the knowledge and capabilities of LLMs and is often followed by rigorous quality filtering or refinement. (1) Style rephrasing techniques transform low-quality, noisy corpora into higher-density expressions. Rephrased web data demonstrates superior utility by aligning with downstream stylistic diversity, such as QA and dialogue, and by enhancing overall data quality (Maini et al., 2024; Akter et al., 2025a). (2) Diffusion synthesis for specific formats typically involves designing specified prompts and generating massive QA pairs or instruction data based on corpora (Cheng et al., 2024; Chen et al., 2025b). Alternatively, it may leverage intermediaries such as personas to enhance diversity and scale (Ge et al., 2025; Zhang et al., 2025e; d). Similarly, code synthesis utilizes designed prompts to generate code instructions based on open-source code snippets (Luo et al., 2024; Wei et al., 2024), or abstracts code into concepts for instruction construction (sc2-instruct). (3) Furthermore, LLM-driven low-resource language translation and multimodal synthesis—such as VLM-driven PDF parsing for data generation (Yang et al., 2025)—extend the utility of data synthesis. Crucially, the core principles underlying these methods are transferable to diverse data formats, providing systematic inspiration for efficient, high-quality synthetic data pipelines during mid-training.
@@ 3.3 Data Selection
The current bottleneck in mid-training data selection lies in filtering pre-training data into high-quality general datasets, predominantly through synergistic combinations of multiple raters (Wettig et al., 2024; Duan et al., 2025; Xu et al., 2025) to cherry-pick optimal samples. Conversely, synthetic or structurally specialized data benefits from more targeted filtering strategies (Zhou et al., 2024), which enable the efficient extraction of domain-aligned, premium content.
@@ 3.5 Data Mixture
Although the determination of optimal data mixing ratios in mid-training inherently depends on the pre-trained model’s capabilities and target objectives, the methodologies for identifying these ratios exhibit broad applicability across different scenarios. Empirically driven allocations remain prevalent; for example, Llama-3 assigns a 30% weight to novel datasets while reserving 70% for its default pre-training data blend (Grattafiori et al., 2024). Beyond heuristic approaches, recent techniques systematically correlate mixing proportions with downstream performance. OLMo 2 (OLMo et al., 2025) employs microanneals, which abbreviate annealing runs on small subsets to extrapolate the efficacy of large-scale mixtures from small-scale trials. Concurrently, RegMix (Liu et al., 2025) formalizes ratio optimization as a regression task, training many small models across diverse mixing configurations to predict unseen mixture performance via regression modeling. These methods hinge on the stability hypothesis, which asserts that ratio performance rankings persist across model scales and token volumes, thereby enabling computationally efficient generalization from proxy experiments to production-level training pipelines.
@@ 8.1 General Capabilities
The preservation and enhancement of general understanding and generation capabilities during mid-training represent a pivotal optimization direction. It requires carefully designed interventions to mitigate catastrophic forgetting while simultaneously refining the model’s specified capabilities. As shown in Table 4, all the mainstream models dive deep into maintaining general performance while improving other objectives. Empirical analyses substantiate that retaining rigorously curated subsets of pre-training data is paramount, optimized composition and proportional balancing of which demonstrably elevate downstream task performance and generalization robustness (Team et al., 2025b). Methodologically, this manifests either through calibrated downsampling of general corpora to amplify specialized data proportions (Gunter et al., 2024; Huo et al., 2025) or via further quality filtration of pre-training sources using advanced classifiers (Parmar et al., 2024; OLMo et al., 2025; Team et al., 2025b). Collectively, these techniques ensure data continuity in the training process while leaving training space for specific data samples.
@@ 8.2.2 Reasoning Capabilities
The systematic enhancement of reasoning capabilities represents a pivotal focus during the mid-training stage, as shown in Table 4, targeting not only the direct improvement of core cognitive dimensions—such as logical deduction, causal inference, and multi-step problem solving—but also establishing foundational groundwork for subsequent reinforcement learning (RL) development (Wang et al., 2025). To achieve this dual objective, prevailing methodologies center on the strategic enrichment of reasoning-intensive data by synthesizing high-quality reasoning samples at scale, including CoT sequences, mathematical proof chains, and structured decision trajectories. These approaches also dynamically calibrate data mixture ratios to progressively elevate the weight of reasoning data without compromising general capabilities (Wake et al., 2025; OLMo et al., 2025; Zhang et al., 2025c). Such data interventions synergize with innovative training strategies, including long-context extension mechanisms that scaffold complex reasoning chains (Hu et al., 2024). As synthetic data can embed diverse, learnable reasoning patterns during mid-training (Abdin et al., 2024b), this intrinsic property positions reasoning optimization as the critical nexus bridging foundation models with downstream capabilities.

### v2

@@ 2 Overview
Mid-training occupies a distinct functional position in the training pipeline, situated between general pre-training and specialized post-training stages. As illustrated in Figure 1, traditional multi-stage training methodologies progressively enhance model performance and generalization by adjusting learning rates and data distributions across different phases (Zhang et al., 2025a). The curriculum learning paradigm naturally aligns with the multi-stage pre-training data curation framework, emphasizing the strategic ordering of training data based on criteria such as data quality, difficulty level, and domain mastery (Naïr et al., 2024; Zhang et al., 2025b). Due to its broad research value and potential, mid-training is increasingly being recognized as a distinct stage, separate from multi-stage training, serving not only as a bridge in terms of data distribution (Liu et al., 2025a) but also enabling targeted adjustments to training strategies and model architectures aimed at developing specific capabilities. Within this emerging classification framework, contemporary LLM development designate pre-training as the foundational stage, where models acquire general capabilities by predicting the next token over large-scale corpora. Emerging as an evolutionary bridge, mid-training preserves these foundational capabilities with the same next-token prediction objective, while systematically enhancing specialized skills through curated data blends—often incorporating high-quality domain-specific content and instruction data. Unlike pre-training, mid-training demonstrates steeper performance gains with higher data efficiency, frequently employing multiple sub-phases for different capability upgrades. In contrast, post-training adopts specialized objectives for SFT/RL and task-specific alignment datasets.
@@ 3.2 Data Synthesis
Distillation leverages powerful LLMs to directly generate data through carefully engineered prompts or to distill smaller synthesis models. This approach is heavily reliant on the knowledge and capabilities of LLMs and is often followed by rigorous quality filtering or refinement.
(1) Style rephrasing techniques transform low-quality, noisy corpora into higher-density expressions. Rephrased web data demonstrates superior utility by aligning with downstream stylistic diversity, such as QA, instruction, and dialogue, and by enhancing overall data quality (Maini et al., 2024; Akter et al., 2025a).
(2) Diffusion synthesis for specific formats typically involves designing specified prompts and generating massive QA pairs or instruction data based on corpora (Cheng et al., 2024; Chen et al., 2025b). Alternatively, it may leverage intermediaries such as personas to enhance diversity and scale (Ge et al., 2025). Building upon these insights, BoostQA (Zhang et al., 2025e) identifies deficiencies in model performance related to STEM disciplines and high-difficulty tasks through probe experiments and proposes a novel large-scale diverse synthesis pipeline focused on STEM and high-difficulty data. Furthermore, LinkQA (Zhang et al., 2025d) proposes a novel strategy that leverages path sampling on the knowledge point graph to construct multi-seed QA synthesis. Similarly, code synthesis utilizes designed prompts to generate code instructions based on open-source code snippets (Luo et al., 2024; Wei et al., 2024b), or abstracts code into concepts for instruction construction (Wei et al., 2024a).
(3) LLM-driven low-resource language translation and multimodal synthesis—such as VLM-driven PDF parsing for data generation (Yang et al., 2025)—extend the utility of data synthesis. Crucially, the core principles underlying these methods are transferable to diverse data formats, providing systematic inspiration for efficient, high-quality synthetic data pipelines during mid-training.
@@ 3.3 Data Selection
The current bottleneck in mid-training data selection lies in filtering pre-training data into high-quality general datasets, predominantly through synergistic combinations of multiple raters (Wettig et al., 2024; Duan et al., 2025; Liangyu et al., 2025) to cherry-pick optimal samples. Conversely, synthetic or structurally specialized data benefits from more targeted filtering strategies (Zhou et al., 2024), which enable the efficient extraction of domain-aligned, premium content.
@@ 3.5 Data Mixture
Although the determination of optimal data mixing ratios in mid-training inherently depends on the pre-trained model’s capabilities and target objectives, the methodologies for identifying these ratios exhibit broad applicability across different scenarios. Empirically driven allocations remain prevalent; for example, Llama-3 assigns a 30% weight to novel datasets while reserving 70% for its default pre-training data blend (Grattafiori et al., 2024). Beyond heuristic approaches, recent techniques systematically correlate mixing proportions with downstream performance. OLMo 2 (OLMo et al., 2025) employs microanneals, which abbreviate annealing runs on small subsets to extrapolate the efficacy of large-scale mixtures from small-scale trials. Concurrently, RegMix (Liu et al., 2025b) formalizes ratio optimization as a regression task, training many small models across diverse mixing configurations to predict unseen mixture performance via regression modeling. These methods hinge on the stability hypothesis, which asserts that ratio performance rankings persist across model scales and token volumes, thereby enabling computationally efficient generalization from proxy experiments to production-level training pipelines.
@@ 8.1 General Capabilities
During mid-training, maintaining and enhancing models’ general understanding and generative capabilities emerges as a crucial optimization direction. This process necessitates carefully designed interventions to mitigate catastrophic forgetting, while simultaneously refining the model’s specialized competencies. As illustrated in Table 4, all mainstream models dive deep into maintaining general performance while improving other objectives. Empirical analyses substantiate that retaining rigorously curated subsets of pre-training data is paramount, optimized composition and proportional balancing of which demonstrably elevate downstream task performance and generalization robustness (Team et al., 2025b). Methodologically, this manifests either through calibrated downsampling of general corpora to amplify specialized data proportions (Gunter et al., 2024; Huo et al., 2025) or via further quality filtration of pre-training sources using advanced classifiers (Parmar et al., 2024; OLMo et al., 2025; Team et al., 2025b). For instance, BoostQA (Zhang et al., 2025e) and LinkQA (Zhang et al., 2025d) facilitate rigorous quality stratification of pre-training datasets, using the QuRater model (Wettig et al., 2024) to quantifies knowledge density and the FineWeb-Edu educational classifier (Penedo et al., 2024) to assess educational utility. Texts rated highly in both knowledge density and educational utility are retained. Collectively, these techniques ensure data continuity in the training process while leaving training space for specific data samples.
@@ 8.2.2 Reasoning Capabilities
Systematic enhancement of reasoning capabilities is one of the central focuses during mid-training, as shown in Table 4. It not only targets the direct improvement of core cognitive dimensions—such as logical deduction, causal inference, and multi-step problem solving, but also lays a robust foundation for subsequent RL development (Wang et al., 2025).
To achieve this dual objective, prevailing methodologies strategically enrich reasoning-intensive datasets through large-scale synthesis of high-quality reasoning samples, including CoT sequences, mathematical proof chains, and structured decision trajectories. For instance, CoTP (Zhang et al., 2025c) proposes leveraging data abundant in diverse, high-value reasoning patterns to expand the model’s reasoning potential, selected by a dual-granularity DTW matching algorithm based on reasoning pattern chains and token entropy chains. Some methodologies further dynamically adjust the data mixing ratios, progressively increasing the proportion of reasoning data without compromising general functionality (Wake et al., 2025; OLMo et al., 2025). Such data interventions synergize with innovative training strategies, including long-context extension mechanisms that scaffold complex reasoning chains (Hu et al., 2024). As synthetic data can embed diverse, learnable reasoning patterns during mid-training (Abdin et al., 2024b), this intrinsic property positions reasoning optimization as the critical nexus bridging foundation models with downstream capabilities.